# PillarTrack: Redesigning Pillar-based Transformer Network for Single Object Tracking on Point Clouds

### v1

@@ Abstract
LiDAR-based 3D single object tracking (3D SOT) is a critical issue in robotics and autonomous driving. It aims to obtain accurate 3D BBox from the search area based on similarity or motion. However, existing 3D SOT methods usually follow the point-based pipeline, where the sampling operation inevitably leads to redundant or lost information, resulting in unexpected performance. To address these issues, we propose PillarTrack, a pillar-based 3D single object tracking framework. Firstly, we transform sparse point clouds into dense pillars to preserve the local and global geometrics. Secondly, we introduce a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design to help the feature representation of each pillar. Thirdly, we present an efficient Transformer-based backbone from the perspective of modality differences. Finally, we construct our PillarTrack tracker based above designs. Extensive experiments on the KITTI and nuScenes dataset demonstrate the superiority of our proposed method. Notably, our method achieves state-of-the-art performance on the KITTI and nuScenes dataset and enables real-time tracking speed. We hope our work could encourage the community to rethink existing 3D SOT tracker designs.
@@ I INTRODUCTION
3D Single object tracking (3D SOT) has a diverse applications in autonomous driving and robotics. Given the initial state (appearance and location) of specific target in the first frame, 3D single object tracking aims to estimate its 3D state across subsequent frames. Existing LiDAR-based 3D single object tracking methods [1, 2, 3, 4, 5, 6, 7] generally follow the Siamese paradigm from 2D visual object tracking, aiming to achieve a trade-off between runtime and accuracy. Given the sparse and irregular input point cloud, these methods initially leverage PointNet families [8, 9] to learn point-wise discriminative representation, and then obtain point-wise similarity with a feature aggregation module. Finally, they estimate the state of specific targets based on these similarity features. SC3D [10] stands as the pioneering LiDAR-based 3D Siamese tracker, leveraging an efficient auto-encoder based on PointNet[11] for point-wise feature encoding. Subsequently, approaches such as P2B [2], 3D-SiamRPN[1], BAT [4] and PTT [3, 12] employ PointNet++ [9] to extract more effective point-wise representation and achieve superior tracking performance. However, one common issue with above point-based 3D SOT methods is the need for re-sampling the input points to a fixed number. For instance. P2B [2] requires the input points to be re-sample to 1024 for the search area and 512 for the template point cloud in order to align the network input. This re-sample operation inevitably introduces the potential for redundant or lost information, which can adversely affect performance. Additionally, the reliance on point query/retrieval in 3D space (e.g., PointNet++ [9]) for the re-sampling process may not be friendly to efficient hardware implementation.
@@
Figure 1: Comparison with other 3D SOT methods on KITTI dataset. We categorize methods based on the Backbone architecture and report their performance on Success and Precision.
@@
To this end, inspired by the recent advancements in pillar-based 3D object detectors [13, 14, 15] due to their real-time speed and high performance. Therefore, we convert the sparse and irregular point clouds into dense and regular pillar representations. It is noteworthy that the 3D SOT task can be regarded as a specific 3D object detection task within a local search area, with the additional template point prior information. Therefore,it is natural to leverage advanced principles from 3D point cloud detection to improve point cloud tracking tasks. Specifically, pillar-based point cloud representation possesses following advantages:(1) Pillar representation is dense and ordered, facilitating seamless integration with advanced 2D image-based techniques without much modification. (2) The compact nature of the pillar representation reduces computational overhead while maintaining a desirable trade-off between performance and speed. (3) Pillar representation is deployment-friendly, making it highly suitable for resource-limited devices like mobile robots or drones. Particularly,pillar representation naturally aligns with the high demanding real-time requirements of 3D trackers, making it well-suited for tracking tasks.
@@
Currently, few works [16, 17] are dedicated to pillar-based 3D SOT with point clouds. SMAT [16] addresses this gap by transforming sparse 3D point clouds into dense pillars and then utilizing a transformer-based encoder for multi-scale feature extraction and interaction, yielding promising results. However, SMAT directly adopts a simplistic pillar encoding module from PointPillars [18] limits the quality of the pillar representations. Besides, their backbone directly follows the design principle of vision transformers tailored for RGB image, which may be sub-optimal for point cloud modalities. PTTR++ [17] introduces a fusion of point-wise and pillar-wise views to further enhance the performance of PTTR [6]. Nevertheless, PTTR++ still involves a re-sampling operation within its pipeline, which presents paratical deployment challenges.
@@
In this paper, we propose PillarTrack, a pillar-based 3D single-object tracking framework, aiming to improve tracking performance while enhancing the inference speed. Firstly, we transform sparse and unordered point clouds into dense and regular pillar representations to reduce information loss caused by re-sampling operation and propose a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design to help the feature learning of each pillar. Secondly, we argue the design of existing Transformer-based backbone from the perspective of modality differences and propose an modality-aware Transformer-based backbone specifically tailored for point cloud modalities, achieving higher performance with fewer GFLOPs. Finally, we construct our PillarTrack network. Extensive experiments on the KITTI and nuScenes dataset demonstrate the superiority of our proposed method. As shown in Figure  1, our approach achieves state-of-the-art performance on the KITTI and nuScenes dataset, striking a better balance between speed and accuracy, and catering to diverse practical needs. In conclusion, the main contributions of this paper are as follows:
@@
$\bullet$ PE-PFE: A Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design to encode the point coordinate of each pillar with pyramid-type representation, and bring the performance gain without extra computational overhead.
@@
$\bullet$ Modality-aware Transformer-based Backbone: A more suitable for point cloud modalities, aiming to enhance feature representation. This design involves a simple adjustment of computational resources allocated to the front-end of the backbone, allowing for capturing more semantic details of the input point cloud.
@@
$\bullet$ SOTA and Open-source: Experiments on KITTI and nuScenes dataset show our method achieve the state-of-the-art performance. Besides, we will open source our code to the research community in https://github.com/StiphyJay/PillarTrack.
@@ II-A Single Object Tracking on Point Clouds
Benefiting from the nature of LiDAR that are less sensitive to illumination changes and capture accurate distance information, numerous works[10, 19, 20, 1, 2, 4, 21, 22] in LiDAR-based 3D Single Object Tracking (3D SOT) have merged. SC3D[10] leverages the Kalman filter for proposals generation and select the most suitable candidate based on Siamese network[23]. 3D-SiamRPN [1] combines 3D Siamese network with 3D RPN network for object tracking. P2B[2] introduces target-specific features and utilizes VoteNet[24] to estimate target center. Based on the P2B pipeline, BAT[4] incorporated Box-aware information to enhance similarity feature. Besides, V2B[22] introduces Voxel-to-BEV network for pixel-wise regression of 3D targets, M2-Track [21, 25] adopts a motion-centric paradigm to address the issue of 3D SOT and achieves remarkable performance. Inspired by the success of transformers[26], several methods [3, 12, 5, 6, 27, 7, 28, 29] have merged. PTT [3, 12] proposed Point-Track-Transformer module to assign weights to crucial point cloud features. LTTR[5] proposes a Transformer-based feature fusion module to learn semantic context among different region. PTTR[6] employs an attention-based feature matching mechanism to integrate target clues in similarity. STNet[7] develops an iterative coarse-to-fine correlation network for robust correlation learning. CXTrack[29] adopts transformer to extract richer contextual information from consecutive frames, enhancing localization accuracy. Recently, SMAT[16] encodes global similarity based on attention in the BEV to alleviate the impact of point sparsity. However, their backbone still partially inherits the design principles from ResNet[30] tailored for the image modality, which may be sub-optimal for point clouds modality.
@@
Figure 2: The architecture of our PillarTrack network.Given the template and search area, we first use PE-Pillar to extract multi-scale features respectively. Then, MAE computes similarity by attention at each feature scale and then fuses the multi-scale similarity feature.Finally, we apply the detection head on the feature fusion map to localize the target.
II-B Pillar-based Object Detection on Point Clouds
Grid-based point cloud 3D detectors[18, 31, 13, 14, 15] generally transform unordered points into regular pillar/voxel and then utilize 2D/3D CNN to learn representation. VoxelNet [32] stands as a pioneering work that leverages Voxel Feature Extractor (VFE) and 3D CNN for 3D object detection. Subsequently, SECOND [33] employs 3D sparse convolution on non-empty voxels to accelerate training and inference. Different from above approaches, PointPillars [18] innovatively converts point clouds into pillars and leverages 2D CNN to enhance detection speed. After that, CenterPoint[31] proposes a center-based head design, achieving milestone performance. Furthermore, PillarNet[13] improves the accuracy of pillar-based 3D detectors through an ”encoder-neck-head” architecture, achieving superior performance while maintaining real-time speed. Based on PillarNet[13], FastPillars[14] proposes a backbone design suitable for point cloud modality and considers deployment-friendly ability, resulting in new state-of-the-art. The success of these pillar-based 3D detectors greatly motivates us to explore the potential of simple pillar-based representation in 3D SOT tasks.
II-C Vision Transformer
@@
Transformer[26] is first proposed in natural language processing (NLP). Subsequently, owing to its remarkable ability for global modeling, the Transformer has shown significant success in image recognition tasks. ViT[34] leverage the Transformer structure to classify images by splitting them into patches. SwinTransformer[35] proposes a shifted windows attention mechanism, achieving excellent performance across various image tasks. PVT[36] and PVTv2[37] utilize a progressive shrinking pyramid to construct a visual Transformer backbone. In the 3D Vison, Point Transformer[38] introduces a specialized Point Transformer layer tailored for processing unordered point clouds. PCT[39] utilizes 3D coordinates of points to generate position embedding and employs an attention module to extract point features from local neighborhoods. These efforts showcase outstanding performance in point cloud shape classification or part segmentation. For LiDAR-based 3D detection tasks, recent approaches[40, 41] first transform point clouds into regular pillars/voxels and then utilize dynamic sparse windows on BEV feature for feature extraction, yielding remarkable performance. In contrast, we employ the Transformer for global modeling on the pillarized feature, aiming to explore the potential of the Transformer in the object tracking task from a BEV perspective.
@@ III Preliminary
III-A Problem Statement
In the task of the 3D Single Object Tracking (3D SOT), we initially define target point cloud (i.e., template) $\mathbf{P_{t}}=\{\mathbf{p}_{ti}=[x_{i},y_{i},z_{i},r_{i}]^{T}\in\mathbb{R}^{N\times 4}\}$, where $x_{i},y_{i},z_{i}$ denote the coordinate values of each point along the axes X, Y, Z, respectively, and $r_{i}$ represents the laser reflection intensity. Our goal is to locate the the 3D bounding box (3D BBox) of this target in the search area $\mathbf{P_{t}}=\{\mathbf{p}_{si}=[x_{j},y_{j},z_{j},r_{j}]^{T}\in\mathbb{R}^{M\times 4}\}$ frame by frame. Here, $N$ and $M$ is the number of points in target and search area. The 3D BBox is represented as $\mathbf{B}=\{\mathbf{b}=[x,y,z,h,w,l,{\theta}]^{T}\in\mathbb{R}^{1\times 7}\}$, The coordinates $x,y,z$ indicate the object’s center, while $h,w,l$ denote its size, ${\theta}_{j}$ is the object’s heading angle. Since the target’s 3D BBox is known in the first frame, we need to regress the target center and heading angle in subsequent frames. By applying the displacement and heading angle from the previous frame’s 3D BBox, we can locate the target in the current frame.
III-B Baseline Architecture Overview
As SMAT represents the first attempt at a pillar-based 3D Siamese tracker and has shown promising performance, we choose SMAT as our baseline and provide a brief overview of its structure. The core components of SMAT include: 1) a naive pillar feature encoding module derived from Pointpillars, which transforms raw point cloud data into a pseudo-image; 2) a Transformer-based backbone network that computes similarity through a multi-head attention mechanism for the multi-scale features from both the template and search branches, capturing local and global information within the point cloud; 3) a neck module that integrates features from different scales and fuses multi-scale similarity features, enhancing the model’s perception of multi-scale information; 4) a decoder that predicts the position and orientation of the target object. These carefully designed features enable SMAT to effectively handle sparse point cloud data, offering an efficient solution for 3D Single Object Tracking.
III-C Motivation
Our research endeavors to investigate a real-time, high-performance, pillar-based approach for 3D single-object tracking. To achieve this, we leverage the compact pillar representation and design a strong Transformer-based backbone that enhances pillar feature representation with reduced computational resources from the perspective of modality differences. Regarding the feature integration and similarity computation module, we follow the original design of baseline method[16], which has shown remarkable efficacy. We hope that our design principles offer a novel perspective in 3D single object tracking, encouraging the community to revisit existing methodologies.
@@ IV Method
IV-A Pyramid-Encoded Pillar Feature Encoding Design
@@
Figure 3: The illustration of PE-PillarVFE design. We encode the channel information of the input point cloud in a pyramid-type fashion. This Pyramid-type encoding design allows the network to optimize more effectively without the input information loss.
@@
In pillar-based 3D detection methods[18, 13, 14], Pillar Feature Encoding (PFE) plays a crucial role, directly impacting the representation learning of BEV feature and final detection accuracy. Its significance extends to 3D single object tracking tasks as well. In SMAT, they utilize the PFE module from PointPillars [18] to extract pillar-wise feature. Here, taking PFE module as an example, we illustrate the details of their input encoding. The points in each pillar are initially concatenated along the channel dimension to form $\hat{p_{i}}=\{[x_{i},y_{i},z_{i},r_{i},{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m},{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]\in{\mathbb{R}^{N_{v}\times 10}}\}$, where $[x_{i},y_{i},z_{i}]$ is the original point coordinates in the ego frame, $[{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m}]$ denotes the offset of $p_{i}$ from the mean of the xyz coordinates of all points within the current pillar, and $[{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]$ is the offset of $p_{i}$ from the current pillar center. However, there exist significant numerical differences in the point-wise feature. For example, the numerical range of the $[x_{i},y_{i},z_{i}]$ is $[-3.2,3.2]$ (i.e., the range of search area). However, the numerical range of $[{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m}]$ and $[{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]$ is $[-0.1,0.1]$, since they are relative offset within a pillar. It can be observed that there exist significant numerical differences (almost 30 $\times$) between the point-wise feature channels in PFE. Such huge disparities in input values pose significant challenges for network optimization  [42, 43]. Therefore, we propose a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design aimed at reducing the numerical differences between the input channels, enabling the network towards better optimization on numerically stable inputs.
TABLE I: Tracking performance comparison of different PFE designs in the Car category in KITTI.
PFE Method Success Precision
PFE (SMAT [16] [^4] The model was trained on a setup comprising four NVIDIA GeForce RTX 3090 GPUs) 68.9 79.6
PE-PFE(Ours) 70.8 (+1.9) 80.3 (+0.7)
Figure 4: The loss comparison of different PFE designs.
As shown in Figure 3, we encode the channel values of the input point cloud in a pyramid-type manner, and finally retain a residual value that cannot be evenly divided. The encoding values from each layer and the residual value are then cascaded together as point-wise features, which are then processed by the PFE module. It is important to note that this pyramid-type encoding design significantly reduces the numerical differences between the input data channels without sacrificing the information contained in the point cloud. This design helps the network achieve better optimization results. Furthermore, to demonstrate the effectiveness of our PE-PFE design compared to the naive PFE from PointPillars [18], we present the performance comparison in the Car category. As shown in Table I, our simple yet effective PE-PFE design yields 1.9 point gains in Success and 0.7 point improvement in Precision. As shown in Figure 4, the design of PE-PFE leads to lower loss values, enabling the network to find better parameters and achieve superior performance.

### v2

# PillarTrack: Boosting Pillar Representation for Transformer-based 3D Single Object Tracking on Point Clouds

@@ Abstract
LiDAR-based 3D single object tracking (3D SOT) is a critical issue in robotics and autonomous driving. Existing 3D SOT methods typically adhere to a point-based processing pipeline, wherein the re-sampling operation invariably leads to either redundant or missing information, thereby impacting performance. To address these issues, we propose PillarTrack, a novel pillar-based 3D SOT framework. First, we transform sparse point clouds into dense pillars to preserve the local and global geometrics. Second, we propose a Pyramid-Encoded Pillar Feature Encoder (PE-PFE) design to enhance the robustness of pillar feature for translation/rotation/scale. Third, we present an efficient Transformer-based backbone from the perspective of modality differences. Finally, we construct our PillarTrack based above designs. Extensive experiments show that our method achieves comparable performance on the KITTI and NuScenes datasets, significantly enhancing the performance of the baseline.
@@ I INTRODUCTION
3D single object tracking (3D SOT) has a wide applications in robotics, such as a robot needs to track their master to complete material handling tasks in factories. Given the initial state (size and location) of target in the first frame, 3D single object tracking aims to estimate its 3D state across subsequent frames. Existing LiDAR-based 3D single object tracking methods [1, 2, 3, 4, 5, 6, 7] generally follow the Siamese paradigm from 2D visual object tracking [8, 9], aiming to achieve a trade-off between runtime and accuracy. Given the sparse and irregular input point cloud, these methods initially leverage PointNet families [10, 11] to learn point-wise representation, and then obtain point-wise similarity with a similarity module. Finally, they estimate the state of specific targets based on these similarity features. SC3D [12] stands as the pioneering LiDAR-based 3D Siamese tracker, leveraging an efficient auto-encoder based on PointNet[10] for point-wise feature encoding. Subsequently, approaches such as P2B [2], 3D-SiamRPN[1], BAT [4] and PTT [3, 13] employ PointNet++ [11] to extract point-wise representation and achieve superior tracking performance. However, one common issue with above point-based 3D SOT methods is the requirement for re-sampling the input points to a fixed number. For instance. P2B [2] re-samples the input points from 1024 to 512. This re-sample operation inevitably introduces the potential for redundant or lost information, which can adversely affect performance.
@@
Figure 1: Comparison with other 3D SOT methods on KITTI dataset. We classify methods based on the backbone architecture and report performance on Success and Precision.
@@
To this end, inspired by the recent advancements in pillar-based 3D object detectors [14, 15] due to their real-time speed and high performance. Therefore, we convert the sparse and irregular point clouds into dense and regular pillar representations. Specifically, pillar-based point cloud representation possesses following advantages: (1) Pillar representation is dense and ordered, facilitating seamless integration with advanced 2D image-based techniques without much modification. (2) The compact nature of the pillar representation reduces computational overhead while maintaining a desirable trade-off between performance and speed. (3) Pillar representation is deployment-friendly, making it highly suitable for resource-limited devices like mobile robots or drones. Particularly, pillar representation naturally aligns with the high demanding real-time requirements of 3D trackers, making it well-suited for tracking tasks.
@@
Currently, there are few works [16, 17] are dedicated to pillar-based 3D SOT tasks. SMAT [16] addresses this gap by transforming sparse 3D point clouds into dense pillars and then utilizing a transformer-based encoder for multi-scale feature extraction and interaction, yielding promising results. However, SMAT utilizes a naive pillar encoding module derived from PointPillars [18], which retains the substantial numerical disparities between point cloud channels. This approach doesn’t fully leverage the intrinsic structure of the point cloud data and limit the representation robustness of pillar. Furthermore, its backbone directly adheres to the design principles of visual transformers [19, 20] that are tailored for RGB images, which is not optimal for the point cloud modality. PTTR++ [17] introduces point-wise and pillar-wise view fusion to further enhance the performance of PTTR [6]. Nonetheless, PTTR++ still incorporates a re-sampling operation in its process, introducing additional computational overhead and posing challenges for practical deployment.
@@
In this paper, we propose PillarTrack, a pillar-based 3D single-object tracking framework, with the objective of augmenting tracking performance and accelerating inference speed, thereby facilitating deployment. Firstly, to reduce information loss caused by re-sampling operation, we propose a Pyramid Encoding Pillar Feature Encoder (PE-PFE) design to transform sparse point clouds into dense and robust pillar representations. Secondly, we argue the design of existing Transformer-based backbone from the perspective of modality differences and propose an modality-aware Transformer-based backbone specifically tailored for point cloud modalities, achieving higher performance with fewer computation resource. Finally, we construct our PillarTrack based on above designs. Extensive experiments on the KITTI and nuScenes dataset demonstrate the superiority of our method. As shown in Figure 1, PillarTrack achieves comparable performance on the KITTI and nuScenes dataset, striking a better balance between speed and accuracy, and catering to diverse practical needs. In conclusion, the main contributions of this paper are as follows:
@@
$\bullet$ PE-PFE: A Pyramid-Encoded Pillar Feature Encoder (PE-PFE) is designed to encode the point coordinate of each pillar with pyramid-type representation. This approach not only enhances performance without incurring additional computational overhead but also improves robustness against rotation, translation, and scale variations.
@@
$\bullet$ Modality-aware Transformer-based Backbone: A tailored for point cloud modalities backbone design, aiming to enhance pillar representation. It involves an adjustment of computational resources allocated to the front-end of the backbone, enabling the capture of more inherent semantic details from the input point cloud.
@@
$\bullet$ Comparable and Open-source: Extensive experiments on KITTI and nuScenes dataset shown the comparable performance, significantly improving upon the baseline approaches. Besides, we open source our code to the community in https://github.com/StiphyJay/PillarTrack.
@@ II Related Work
Single Object Tracking on Point Clouds. Leveraging the advantages of LiDAR, which is less sensitive to illumination changes and captures accurate distance information, numerous works[12, 21, 1, 2, 4, 22, 23] in LiDAR-based 3D Single Object Tracking (3D SOT) have merged. SC3D[12] utilizes a Kalman filter for proposals generation and selects the most suitable candidate via a Siamese network. 3D-SiamRPN [1] combines 3D Siamese network with 3D RPN network for tracking. P2B[2] introduces target-specific features and employs VoteNet[24] to estimate the target center. Based on this, BAT[4] incorporates box-aware information to enhance similarity feature. M2-Track [22, 25] adopts a motion-centric paradigm that yields impressive results. Inspired by the success of transformers[26], several transformer-based methods [3, 13, 5, 6, 27, 7, 28, 29] have merged. PTT [3, 13] proposed Point-Track-Transformer module to assign weights to crucial point cloud features. PTTR[6] employs an attention-based feature matching mechanism to integrate target clues in similarity. STNet[7] develops an iterative coarse-to-fine correlation network for robust correlation learning. GLT-T[28] proposes a global-local Transformer voting scheme to generate higher quality 3D proposals. Recently, SMAT[16] encodes global similarity based on attention to alleviate the point sparsity. However, their backbone still partially inherits the design principles from ResNet[30] tailored for the image modality, which is sub-optimal for point clouds. Therefore, we argue their design and propose a modality-aware transformer-based backbone tailored for point cloud modality.
@@
Figure 2: The architecture of our PillarTrack network. Given the template and search area, we first use PE-PFE to extract multi-scale features respectively. Then, the backbone extract features at different feature scale and fuses the multi-scale similarity feature. Finally, we apply the detection head on the feature fusion map to localize the target.
@@
Vision Transformer. The Transformer architecture [26], originally proposed for natural language processing (NLP), has demonstrated significant success in image recognition due to its global modeling capabilities. The Vision Transformer (ViT)[31] classifies images by dividing them into patches, while Swin Transformer[32] introduces a shifted window attention mechanism, achieving excellent performance across various visual tasks. PVT [19] and PVTv2 [20] employ a progressive shrinking pyramid to construct a visual Transformer backbone. In 3D vision, Point Transformer [33] features a specialized Point Transformer layer for processing unordered point clouds. These methods have shown outstanding performance in point cloud shape classification and part segmentation. In contrast, our approach explores the potential of the Transformer’s global modeling on pillar features in the object tracking task.
@@ III Preliminary
Motivation. Our research endeavors to investigate a real-time, high-performance, pillar-based approach for 3D single-object tracking. To achieve this, we leverage the compact pillar representation and design a strong Transformer-based backbone that enhances pillar feature representation with reduced computational resources from the perspective of inherent modality differences. Regarding the feature integration and similarity computation module, we follow the original design of baseline[16], which has shown remarkable efficacy. We hope that our designs can offer a novel perspective in 3D SOT, encouraging the community to revisit existing transformer-based methods.
Task Definition. In the task of the 3D Single Object Tracking (3D SOT), we initially define target point cloud (i.e., template) $\mathbf{P_{t}}=\{\mathbf{p}_{i}^{t}=[x_{i},y_{i},z_{i},r_{i}]^{T}\in\mathbb{R}^{N\times 4}\}$, where $x_{i},y_{i},z_{i}$ denote the coordinate values of each point along the axes X, Y, Z, respectively, and $r_{i}$ represents the laser reflection intensity. Our goal is to locate the the 3D bounding box (3D BBox) of this target in the search area $\mathbf{P_{t}}=\{\mathbf{p}_{si}=[x_{j},y_{j},z_{j},r_{j}]^{T}\in\mathbb{R}^{M\times 4}\}$ frame by frame. Here, $N$ and $M$ is the number of points in target and search area. The 3D BBox is represented as $\mathbf{B}=\{\mathbf{b}=[x,y,z,h,w,l,{\theta}]^{T}\in\mathbb{R}^{1\times 7}\}$, The coordinates $x,y,z$ indicate the object’s center, while $h,w,l$ denote its size, ${\theta}_{j}$ is the object’s heading angle. Since the target’s 3D BBox is known in the first frame, we need to regress the target center and heading angle in subsequent frames. By applying the displacement and heading angle from the previous frame’s 3D BBox, we can locate the target in the current frame.
Baseline Overview. As SMAT is the first attempt at a pillar-based 3D Siamese tracker and has shown promising results, we choose SMAT as our baseline and briefly review its structure. The core components of SMAT include: 1) a naive pillar feature encoding module derived from Pointpillars, which transforms raw point cloud data into a pseudo-image; 2) a Transformer-based backbone network that computes similarity through a multi-head attention mechanism for the multi-scale features from both the template and search branches, capturing local and global information within the point cloud; 3) a neck module that integrates features from different scales and fuses multi-scale similarity features, enhancing the model’s perception of multi-scale information; 4) a decoder that predicts the position and orientation of the target object. These designs enable SMAT to effectively handle sparse point cloud, offering an efficient solution for 3D SOT.
@@ IV Method
IV-A Pyramid-Encoded Pillar Feature Encoding
@@
Figure 3: The illustration of PE-PFE design. We encode the point coordinate of the input point cloud in a pyramid-like type. This Pyramid-like encoding design allows the network to optimize effectively without the input information loss.
Figure 4: Illustration of the properties of PE-PFE. The red indicates the feature distribution after original PFE encoding, while the blue represents the feature distribution after PE-PFE encoding. (a) Original point cloud. (b) Point cloud with 1.2X scale. (c) Point cloud with 1.2m translation. (d) Point cloud with 45°$\degree$ rotation.
Figure 5: The output features visualization of forth stage in t-SNE.(a) Visualization with the image blocks setting. (b) Visualization matching the total number of the original blocks. (c) Visualization after removing redundancies from the number of blocks.
@@
In pillar-based 3D detection methods[18, 14, 15], Pillar Feature Encoding (PFE) plays a crucial role, directly impacting the representation learning of BEV feature and final detection accuracy. Its significance also extends to 3D single object tracking tasks. SMAT utilizes the PFE from PointPillars [18] to extract pillar-wise feature. Here, taking PFE module as an example, we illustrate the details of their input encoding. The points in each pillar are initially concatenated along the channel dimension to form $\hat{p_{i}}=\{[x_{i},y_{i},z_{i},r_{i},{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m},{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]\in{\mathbb{R}^{N_{v}\times 10}}\}$, where $[x_{i},y_{i},z_{i}]$ is the original point coordinates in the ego frame, $N_{v}$ is the number of points within the pillar, $[{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m}]$ denotes the offset of $p_{i}$ from the mean of the xyz coordinates of all points within the current pillar, and $[{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]$ is the offset of $p_{i}$ from the current pillar center. However, there exist large numerical differences between the feature channels in PFE. For instance, the numerical range of $[{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m}]$ and $[{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]$ is $[-0.1,0.1]$, since they are relative offset within a pillar. Such huge disparities in input values pose significant challenges for network optimization [34] and can make model sensitive to disturbance, such as scaling, translation or rotation. To this end, we propose the Pyramid-Encoding Pillar Feature Encoder (PE-PFE), which is designed to minimize value discrepancies between channels, thereby facilitating improved optimization of numerically stable inputs. Besides, PE-PFE ensures that input point clouds maintain robustness against scale, translation, rotation.
In detail, as shown in Figure 3, we encode the channel values of the input point cloud in a pyramid-type manner, and finally retain a residual value that cannot be evenly divided. The encoding values from each layer and the residual value are then cascaded together as point-wise features, which are then processed by the PFE module. Similar to the natural language processing (NLP) tokenization, the PE-PFE serves as a feature embedding that converts point cloud data into an information-dense representation. The embedding of PE-PFE is characterized by threes properties: scale-robustness, translation-robustness, and rotation-robustness. Scale robustness aids in the feature extraction of geometrical information for vehicles with identical shapes but varying sizes, while remaining insensitive to scale variations. Translation-robustness ensures that even with positional perturbations in the input point cloud, the distribution of feature encoding remains stable, facilitating reliable representation. Rotation-robustness assists in extracting point cloud with different orientation angle, enhancing the expression of semantics. Collectively, these properties contribute to the robustness and effectiveness of PE-PFE in feature extraction and geometric understanding within the 3D SOT tasks. For instance, as show in Figure 4, we apply different operations of 1.2X scale, 1.2m translation, and 45°$\degree$ rotation to the raw point cloud of a vehicle. The input feature distribution of original PFE exhibits significant discrepancies; however, after PE-PFE, the feature distributions remained more stable. We remain the exploration of PE-PFE on other Lidar-based tasks in the future work. For the effectiveness of PE-PFE, refer to Table V.